OpenStudy (anonymous):

Find the point of intersection y=3x+4 y=-2x+14

OpenStudy (anonymous):

set them equal and solve for \(x\) since \(y=y\) you get \[3x+4=-2x+14\]

OpenStudy (anonymous):

i got 3,8

OpenStudy (anonymous):

i don't think that is right, lets check

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

0,2?

OpenStudy (anonymous):

\[3x+4=-2x+14\] add \(2x\) get \[5x+4=14\] subtract 4 get \[5x=10\] divide by 5 get \[x=2\]

OpenStudy (anonymous):

2,10? is the answer

OpenStudy (anonymous):

yes